. You joined a “dogsled” race during your winter break. To start, you

pulled the sled (total mass of 85 kg) with a force of 650N at 50° above

the horizontal. Find the work done on the sled after it moves at a

distance of 800cm. (Round to 2 decimal places)






1
Expert's answer
2021-11-07T19:25:39-0500

"s = 800cm = 8m"

"F = 650N"

"\angle \alpha= 50\degree"

"m = 85 kg"

"W = Fs\cos \alpha= 650*8*\cos 50\degree= 3342.49557\approx3342.50J"


"\text{Answer: }3342.50J"